9/10 volume of an ice cube is floating down the surface of sea water.find the density of the ice cube
1
Expert's answer
2018-10-10T10:28:09-0400

The equilibrium condition between gravity force and buoyancy gives

mg=gρ_W V_disp

Here V_disp=9/10 V is a volume of the displaced body of liquid.

Since

m=ρ_ice V

We get

ρ_ice V=ρ_W 9/10 V

ρ_ice=9/10 ρ_W=9/10×1027 kg/m^3 =924.3 kg/m^3